# Truck Parking Usage Prediction with Decomposed Graph Neural Networks

### v1

@@ III-B Regional Temporal Graph Convolutional Network
Inspired by Attention Temporal Graph Convolutional Network (A3TGCN) [27], we built the Graph Neural Network, which incorporates the geographical relations in nodes integrating spatial and temporal aspects with regional graph inputs. Our model comprises two main components: a Graph Convolution Network (GCN) as a spatial module and a recurrent unit as a temporal module. GCN in the spatial convolution module handles the node’s attributes, and a specialized GCN residing in Regional Decomposition compresses the spatial dependencies between sites by considering the distances between each pair of truck parking sites. Given the sequence of input timestamps $t\in\{1,2,\ldots,K\}$, the temporal model predicts site occupancy rates $o$ up to $T$ future time steps.
For each time step $t$, in the node embedding at a graph level, we consider $\hat{A}$ as the adjacency matrix, wherein $\hat{A}=A+I_{N}$ with $I_{N}$ being the identity matrix. Additionally, $\hat{D}$ represents a diagonal degree matrix, and its elements are defined as $\hat{D}_{ii}=\sum_{j=0}\hat{A}_{ij}$ and the weight matrix is denoted as $W$. It demands $H_{t+1}$ of the hidden vectors. Following the equation proposed by Kipf and Welling [21], this progression can be expressed as:
@@
        $M^{r}_{t-k}\leftarrow\text{RD}(G_{t-k})$
@@
     $X_{t+1}\leftarrow\text{Decode}(H_{t+1})$
@@
To leverage the potential benefits of regional relationships, we introduce the Regional Decomposition technique. This involves the creation of subgraphs $M$ while decomposing the original graph of truck parking site locations $G$. During each training step, the Regional Decomposition process distributes each node attribute $\eta$ by regional relationships. This process leads to deriving independent vectors $\{m,\ldots,m_{N}\}\in\mathnormal{R}^{N\times E}$ with the number of features $E$.
Input graphs are sequentially aligned at every 10-minute interval and passed to spatial modules to extract the node features. Each feature represents regional graph structural information and nodes feature-based information within the regional context. Our whole algorithm is shown in Algorithm 1. Each GCN is applied to the regional graphs to extract spatial features as hidden vectors $m$. Subsequently, these vectors are concatenated, and the linear layer computes the hidden vector $\gamma$, which can be used in the temporal layers. These operations are performed iteratively for each time step.
@@
Formally, we arbitrarily collect the region of the nodes in $G$ and enforce them to connect together as a graph $M_{r}$. On the subgraph networks, we can consider $\gamma$ as a concatenated feature derived by the node embeddings spatial GCN denoted as
$\gamma=\bigcup^{r}_{i=1}(m^{i},e_{r_{i}})\subseteq G$ (8)
@@ III-C Temporal Module
The temporal model is applied to acquire temporal dependencies with hidden vectors from the previous time step. To obtain the temporal dependencies of composed features, we constructed the GRU architecture using GCN, inspired by the baseline model [27]. This is the extension technique of Convolutional LSTM [28] to handle spatial dependencies into a temporal module. Figure 3 illustrates the architecture of GRU, which can accommodate the graph structure and composed features. Within this framework, activation function $\sigma$ represents the sigmoid function, and parameters $W$ correspond to the weights. The function $f(v_{t},A_{t})$ denotes the computation of graph convolutions at time step $t$, which is subsequently concatenated with input hidden vectors. Accordingly, the structural convolution function can be formally defined as follows:
$f(v,A)=\sigma(W\eta_{i}+W\sum^{d_{i}}_{k}\eta_{i}^{(k)})$ (9)
Incorporating the structural formulas and temporal dependency function, our RegT-GCN formulation at the node level can be expressed as follows with the initial condition of $h_{0}=\gamma_{0}$.
@@ III-D Model Training
After iterating input time steps in the temporal module, they are all passed to the decoder, consisting of two linear layers and the ReLU activation function.
We also modified the attention vector $\alpha_{t}$ as an aggregation function with a simplified form to enhance the computational efficiency and in accordance with our experimental decision.
$\alpha_{t}=\frac{\exp(\alpha_{t})}{\sum_{s\in\mathnormal{N}(K)\setminus t}\exp(\alpha_{s})}$ (11)
Finally, the calculated hidden vectors are passed through the decoder to yield the predicted outputs for nodes at time step $t+1$, denoted $X_{t+1}$, and the hidden vector matrix is expressed as $H_{t}$.
$\begin{split}X_{t+1}&=\text{ReLU}(H_{t}W^{(0)}+b^{(0)})W^{(1)}+b^{(1)}\\ &=\text{ReLU}(\text{ReLU}(\alpha_{t}f(X_{t},A_{t})\oplus H_{t-1})W^{(0)}\\ &+b^{(0)})W^{(1)}+b^{(1)}\end{split}$ (12)
During the training process, the model is trained to minimize the discrepancy between the actual truck parking occupancy rates and predicted values across time steps. The model loss is then calculated using the target attribute, occupancy rates matrix $O_{t}$, employing the mean squared error criterion.
$\mathnormal{L}=\frac{1}{n}\sum(O_{t}-\hat{O_{t}})^{2}$ (13)
where $n$ is the total number of parking sites and $\hat{O_{t}}$ is the predicted occupancy rates matrix.
@@ IV Experiments
This section explains data, data handling, other experiment settings, and the results. There are two perspectives of results, including truck parking usage prediction in different time horizons and the efficiency of graph connectivity regarding the model performance of Regional Decomposition. Discussion follows in the end.
@@ IV-A Data Collection and Pre-processing
A collaborative effort among eight states in MAASTO led to the development of a real-time, multi-state Truck Parking Information Management System (TPIMS). Operational since 2019, the TPIMS data archival system (TPIMS-DAS) has archived over 100 million records of truck parking information (until October 2023).
@@
Figure 4 shows all truck parking sites in the MAASTO states and where we have collected parking data. Currently, TPIMS-DAS has archived data from 144 parking sites (public and private) of the eight participating MAASTO states. The number of current sites reporting data in each state is Iowa: 44, Illinois: 19, Kansas: 18, Kentucky: 13, Michigan: 14, Minnesota: 7, Ohio: 18, and Wisconsin: 11. About 160,000 parking records are archived in TPIMS-DAS daily; hence more than 1000 records per site on average. Due to maintenance periods in some sites and a lack of consequent data, we picked 105 sites for the graph construction and usage prediction.
Before applying our model, we handled the raw data in the appropriate format. We extracted dependent features from archived data following the formats in Table I. Week ID, day ID, and hour ID are dynamically changed, and other features are statically decided. These features were identity encoded and used as the input features of nodes. The occupancy rate is calculated using available data and capacity and the target feature of the model.
Edges are composed of two different sites, and with the characteristics of road networks, not all nodes are connected. To realize that, we selected the sites within 40 miles, which is approximately a 30-minute driving distance, and connected them for a single connected graph and regional subgraphs. The randomly connected graph used to train models connects all sites within the same group. Each edge has the actual driving distance as weights extracted from Bing Maps Locations API. One problem exists: data is collected in different frequencies. For this research, we used 10-minute frequency data, but some sites have higher data collection frequencies. Therefore, we interpolated the occupancy rate data, taking the average of two consecutive values to fill in the blank data. Other dynamic features were determined as the same number of previous values.
@@ IV-B Evaluation Metrics
Usage prediction metrics compare predicted value and ground truth data statistics. They are computed using a root mean squared error (RMSE), mean absolute error (MAE), and mean absolute percentage error (MAPE). We focused on the occupancy rates on every site, which is the parking usage divided by the full capacity. The distance between the predicted and ground truth value is interpreted as the prediction accuracy, and RMSE also considers the outliers by providing a large penalty.
@@
IV-C Comparison Models
We compared our proposed framework, Regional Temporal GCN (RegT-GCN), with the following baseline methods: Stacked GRU, Stacked GCN [29], Temporal GCN (T-GCN) [27], Temporal Graph Convolution LSTM (LTGC) [30]. We also built the comparison methods for the ablation study: Convolution Stacked Temporal GCN (CST-GCN), Random Temporal GCN (RanT-GCN), and Regional Temporal GCN with LSTM (RegLT-GCN). They were all applied RMSProp and the learning rate is $10^{-3}$ with the decaying rate of $10^{-4}$. Input channels are 8 for all models, and output channels are 512 for LSTM, T-GCN, and LTGC and 256 for the other models.
@@ IV-C Comparison Models
  • T-GCN is composed of GRU as a temporal module with attention based aggregation named Attention Temporal GCN (A3T-GCN) on its spatial module [27].
@@
  • CST-GCN applies multiple convolution layers sequentially to input nodes and compares the performance of the vector concatenation. Each calculated hidden vector is inputted to the next convolution layers, a total of five layers, the same as five convolutions for regions in RanT-GCN and RegT-GCN.
  • RanT-GCN is analogous to DropEdge [31], which applied random edge sampling on the graph to make several subgraphs. Most modules are the same as RegT-GCN but use random connected graphs as input subgraphs.
@@ IV-D Truck Parking Usage Prediction
Ours
CST-GCN 0.150 0.179 0.183 0.199 0.117 0.140 0.142 0.156 16.56 19.84 19.65 21.04
RanT-GCN 0.089 0.095 0.119 0.151 0.064 0.068 0.088 0.115 9.12 9.57 12.39 15.93
RegT-GCN 0.086 0.094 0.120 0.150 0.062 0.067 0.088 0.114 8.76 9.36 12.43 15.68
Table II demonstrates the performance comparison among different models. Firstly, Stacked GCN predicted better than Stacked GRU, which indicates the spatial model learns better representations of truck parking sites than the temporal model. Longitudinally, Stacked GRU lowered its performance than Stacked GCN since the temporal model highly depends on the predicted time horizons. Next, most spatio-temporal models resulted in better accuracy, especially in shorter time horizons compared to Stacked GRU and Stacked GCN. At longer time horizons, the result of Stacked GCN predicted better or closer accuracy to spatio-temporal models, while Stacked GRU and T-GCN performance degraded the prediction accuracy.
In the ablation study of the Regional Decomposition module, RanT-GCN and RegT-GCN surpassed the baseline models in most of the prediction time horizons. Within these two, decomposing a graph regionally achieved the best accuracy of all. CST-GCN predicted the lowest accuracy; hence, the stack of convolution layers doesn’t prove the model’s superiority over the models integrated with the decomposition technique.
Figure 5 (a) and (b) illustrate the comparison of the models trained on connected graphs, randomly decomposed graphs, and regionally decomposed graphs on one specific truck parking site. In the graph, $1.0$ indicates that the parking lots are fully occupied. In the figure, T-GCN predicted higher occupancy rates in both observations. The prediction results of RegT-GCN and RanT-GCN are analogous to the ground truth results, which also adequately predicted the intense usages. Meanwhile, RanT-GCN predicted the values in tendencies of fewer fluctuations than RegT-GCN. There is no significant difference in observed patterns between the results of the original model and the model generalization.
Similarly, Figure 5 (c) and (d) show the prediction results of future truck parking usage for an adequately long time horizon (6 hours). In the figure, the accuracy of each model is lower than the outcomes of a time horizon of 30. Specifically, T-GCN with a connected graph has the lowest prediction capability in both results and has tendencies to predict lower occupancy rates. Both RanT-GCN and RegT-GCN captured the usage characteristics except for the intense usage, which proves that graph decomposition can also be effective in longer horizons such that the truck drivers can schedule their route before their working time. Moreover, RegT-GCN predicted closer occupancy rates to the ground truth.
IV-E Models Generality
Table III: The generality comparison of prediction results using baseline and our models on the truck parking dataset with different time horizons (displayed in minutes).
Models/Metrics Time Horizons (10/30/120/360)
@@ IV-E Models Generality
To assess the generalizability of the models, we selected additional weeks for the test sequences and conducted inference experiments with the pre-trained models. Table III summarizes the results of RMSE, MAE, and MAPE to evaluate the model’s predictive accuracy. In the table, every model generalizes its predictive abilities to other locations except for CST-GCN.
@@
Further analysis reveals that the RanT-GCN and RegT-GCN consistently outperformed the baseline models across various time horizons, indicating their superior predictive capabilities. Remarkably, even when baseline models exhibit lower performance than the original prediction results, models with integrated decomposition techniques maintain their efficacy in providing accurate and generalizable inferences. These findings underscore the consistent and superior performance of the Regional Decomposition in predictive accuracy, regardless of the time horizon under consideration.
Figure 5: Results on Regional Decomposition for the time horizon of 30 (a and b) and 360 (c and d).
IV-F Discussion

### v2

@@ III-B Regional Temporal Graph Convolutional Network
In this paper, we built the Graph Neural Network, which incorporates the geographical characteristics in nodes integrating spatial and temporal aspects with regional graph input. Our model comprises two main components: a Graph Convolution Network (GCN), inspired by the Attention Temporal Graph Convolutional Network (A3TGCN) [27], as a spatial module and a recurrent unit as a temporal module. A structual GCN handles the node embeddings of the truck parking location graph, and a spatial GCN compresses the spatial dependencies between sites taken the distances between each pair of truck parking sites. Given the sequence of input timestamps $t\in\{1,2,\ldots,K\}$, the temporal model predicts site occupancy rates up to $T$ future time steps.
For each time step $t$, in the node embedding at a graph level, we denote $\hat{A}$ as the adjacency matrix with self-loop, where $\hat{A}=A+I$ with $I$ being the identity matrix. Additionally, $\hat{D}$ represents a diagonal degree matrix, its elements are defined as $\hat{D}_{ii}=\sum_{j=0}\hat{A}_{ij}$ and the weight matrix is denoted as $W$. It demands $H_{t+1}$ of the hidden vectors. Following the equation proposed by Kipf and Welling [21], this progression can be expressed as:
@@
        $M_{t-k}\leftarrow\text{Regional Decomposition}(G_{t-k})$
@@
     $V_{t+1}\leftarrow\text{Decode}(H_{t+1})$
@@
To leverage the potential benefits of regional relationships, we introduce the Regional Decomposition technique. This involves the creation of subgraphs $M$ while decomposing the original graph of the truck parking site locations $G$. During each training step, the Regional Decomposition process distributes each node attribute $\eta$ by regional relationships. This process leads to deriving independent vectors $\{m_{1},\ldots,m_{N}\}\in\mathcal{R}^{N\times E}$ with the number of features $E$ and the total number of truck parking site $N$.
The input graphs are sequentially aligned at every 10-minute interval and passed to spatial modules to extract the node features. Each feature represents structural information from the regional graph and information from the nodes based on features within the regional context. Our whole algorithm is shown in Algorithm 1. Each GCN is applied to the regional graphs to extract spatial features as hidden vectors $m$. Subsequently, these vectors are concatenated, and the linear layer computes the hidden vector $\gamma$, which can be used in the temporal layers. These operations are performed iteratively for each time step.
@@
Formally, we arbitrarily collect the region of the nodes in $G$ and enforce them to connect together as a graph $M^{r}$. On the subgraph networks, we can consider $\gamma$ as a concatenated feature derived by the node embeddings from spatial GCN denoted as
$\gamma=\bigcup_{r}(m^{r},e^{r})$ (8)
@@ III-C Temporal Module
The temporal model is applied to acquire temporal dependencies with hidden vectors from the previous time step. To obtain the temporal dependencies of composed features, we constructed the GRU architecture using GCN, inspired by the baseline model [27]. This is the extension technique of Convolutional LSTM [28] to handle spatial dependencies into a temporal module. Figure 2 illustrates the architecture of GRU, which can accommodate the graph structure and the composed features. Within this framework, the activation function $\sigma$ represents the sigmoid function, and the parameters $W$ correspond to the weights. The function $f(v_{t},A_{t})$ denotes the computation of graph convolutions in the time step $t$, which is subsequently concatenated with the input hidden vectors. Accordingly, the structural convolution function can be formally defined as follows:
$f(v,A)=\sigma(W\gamma_{i}+W\sum^{d_{i}}_{l}\gamma_{i}^{(l)})$ (9)
Incorporating the structural formulas and the temporal dependency function, our RegT-GCN formulation at the node level can be expressed as follows with the initial condition of $h_{0}=\gamma_{0}$.
@@ III-D Model Training
After iterating the input time steps in the temporal module, they are all passed to the decoder, consisting of two linear layers and the ReLU activation function.
We also modified the attention vector $\alpha_{t}$ as an aggregation function in simplified form to improve computational efficiency and in accordance with our experimental decision.
$\alpha_{t}=\frac{\exp(\alpha_{t})}{\sum_{s\in\mathcal{N}(K)\setminus t}\exp(\alpha_{s})}$ (11)
Finally, the calculated hidden vectors are passed through the decoder to yield the predicted outputs for the nodes at the time step $t+1$, denoted $V_{t+1}$, and the hidden vector matrix is expressed as $H_{t}$.
$\begin{split}V_{t+1}&=\text{ReLU}(H_{t}W^{(0)}+b^{(0)})W^{(1)}+b^{(1)}\\ &=\text{ReLU}(\text{ReLU}(\alpha_{t}f(V_{t},A_{t})\oplus H_{t-1})W^{(0)}\\ &+b^{(0)})W^{(1)}+b^{(1)}\end{split}$ (12)
During the training process, the model is trained to minimize the discrepancy between the actual truck parking occupancy rates and predicted values across time steps. The loss of the model is then calculated using the target attribute, the occupancy rates matrix $O_{t}$, using the mean squared error criterion.
$\mathcal{L}=\frac{1}{N}\sum(O_{t}-\hat{O_{t}})^{2}$ (13)
where $\hat{O_{t}}$ is the matrices of predicted occupancy rates.
@@ IV Experiments
This section explains data, data handling, experiment settings, and the results. There are two perspectives of results, including truck parking usage prediction in different time horizons and the efficiency of graph connectivity regarding the model performance of Regional Decomposition. Discussion follows in the end.
@@ IV-A Data Collection and Pre-processing
A collaborative effort among eight states in MAASTO led to the development of a real-time multi-state Truck Parking Information Management System (TPIMS). Operational since 2019, the TPIMS data archival system (TPIMS-DAS) has archived more than 100 million records of truck parking information (up to October 2023).
@@
Figure 3 shows all truck parking sites in MAASTO states and where we have collected parking data. Currently, TPIMS-DAS has archived data from 144 parking sites (public and private) of the eight participating MAASTO states. The number of current sites reporting data in each state is Iowa: 44, Illinois: 19, Kansas: 18, Kentucky: 13, Michigan: 14, Minnesota: 7, Ohio: 18, and Wisconsin: 11. About 160,000 parking records are archived in TPIMS-DAS daily; hence more than 1000 records per site on average. Due to maintenance periods in some sites and a lack of consequent data, we picked 105 sites for the graph construction and usage prediction.
Before applying our model, we handled the raw data in the appropriate format. We extracted dependent features from archived data following the formats in Table I. Week ID, day ID, and the hour ID are dynamically changed, and other features are statically decided. These features were identity encoded and used as the input features of the nodes. The occupancy rate is calculated using the available data and capacity, and is the target feature of the model.
Edges are composed of two different sites, and with the characteristics of road networks, not all nodes are connected. To realize that, we selected sites within 40 miles, which is approximately a 35-minute driving distance, and connected them for a single connected graph and regional subgraphs. Each edge has the actual driving distance as weights extracted from Bing Maps Locations API. For this research, we used 10-minute frequency data, while some sites have different data collection frequencies. Therefore, we interpolated the occupancy rate data, taking the average of two consecutive values to fill in the blank data. Other dynamic features were determined to be the same number as previous values.
@@ IV-B Evaluation Metrics
Usage prediction metrics compare the predicted value and ground truth data statistics. They are calculated using the root mean squared error (RMSE), mean absolute error (MAE), and mean absolute percentage error (MAPE). We used the occupancy rates on every site for the target value, which is the parking usage divided by the full capacity. The distance between the predicted and ground truth values is interpreted as the prediction accuracy. In addition, RMSE also considers the outliers by providing a large penalty.
@@
IV-C Model Comparison
We compared our proposed framework, Regional Temporal GCN (RegT-GCN), with the following baseline methods: Stacked GRU, Stacked GCN [29], Temporal GCN (T-GCN) [27], Temporal Graph Convolution LSTM (LTGC) [30] and Graph SAGE [31]. They were all applied RMSProp and the learning rate is $10^{-3}$ with the decaying rate of $10^{-4}$. Input channels are 8 for all models and output channels are 512 for LSTM, T-GCN, and LTGC and 256 for the other models. For this experiment, we used one GeForce RTX 3080 for training and inference. Experimentally, our model takes around 15 minutes for training and 1 minute for inference.
@@ IV-C Model Comparison
  • T-GCN is composed of GRU as a temporal module with attention-based aggregation named Attention Temporal GCN (A3T-GCN) on its spatial module [27].
@@
  • Graph SAGE is a general framework for inductive representation learning on large graphs with low-dimensional node embeddings.
@@ IV-D Truck Parking Usage Prediction
RegT-GCN (ours) 0.086 0.094 0.120 0.150 0.062 0.067 0.088 0.114 8.76 9.36 12.43 15.68
Figure 4: The results of occupancy rate prediction with RegT-GCN at one truck parking site for the time horizon of 30 minutes (above) and 360 minutes (below) and ground truth from Mar. 3rd, 2022 to Mar. 14th, 2022.
Each model was trained on $20\%$ of the data (March 1st to March 3rd, 2022) and tested on the rest of $80\%$ of the data (March 3rd to March 14th, 20202). Table II demonstrates the performance comparison with baseline models. Comparing the temporal and spatial models, Stacked GCN predicted generally better than Stacked GRU, which indicates the spatial model learns better representations of truck parking sites than the temporal model. In comparison, most spatio-temporal models improved the model performance, especially in shorter time horizons compared to Stacked GRU and Stacked GCN. Across the models, our RegT-GCN outperforms the baselines. Specifically in longitudinal analysis, Stacked GRU decreased its performance compared to Stacked GCN since the temporal model highly depends on the predicted time horizons. At longer time horizons, the result of Stacked GCN predicted a better or closer accuracy to spatio-temporal models, while Stacked GRU and T-GCN performance dropped the prediction accuracy. Across most time horizons, our RegT-GCN successfully demonstrates improvements in truck parking prediction performance.
Figure 4 illustrates the predicted results of RegT-GCN and ground truth. The above figure compares the predicted occupancy rate over the timestamp beginning from Mar. 3rd, 2022. The prediction results show that the model successfully learns the tendencies of truck parking usage, which also adequately predicts intense usage. The figure below shows the predicted occupancy rate over a time horizon of 360 minutes. RegT-GCN also learns the tendency of the usage at whole timestamps. However, compared to a time horizon of 30 minutes, RegT-GCN exhibits limitations in predictive performance and capturing peak values accurately. These examples prove that graph decomposition can be effective in any time horizon, such that truck drivers can schedule their route anytime.
IV-E Model Generality
Table III: The generalizability comparison of prediction results using baseline and our models on the truck parking dataset with different time horizons (displayed in minutes).
@@ IV-E Model Generality
To assess the generalizability of the models, we selected additional weeks, starting from March 15th, 2022, for the test sequences and conducted inference experiments with the pretrained models. Table III summarizes the results of RMSE, MAE, and MAPE to evaluate the model’s predictive accuracy. In the table, every model shows the general capability of its predictive abilities to other time ranges.
@@
Further analysis reveals that RegT-GCN consistently outperformed baseline models across various time horizons, indicating their superior predictive capabilities. Remarkably, even when baseline models exhibit lower performance than the original prediction results, models with integrated decomposition techniques maintain their efficacy in providing accurate and generalizable inferences. These findings underscore the consistent and superior performance of the Regional Decomposition in predictive accuracy, regardless of the time horizon under consideration.
IV-F Regional vs Random Graph Decomposition
We also conducted an experiment on graph connectivity exploration. Figure 5 illustrates the diversity of connection approaches. We divided the truck parking location graph into subgraphs characterized by random connections. The proposed model, Random T-GCN (RanT-GCN), is analogous to DropEdge [32], which applied random edge sampling on the graph to make several subgraphs. Most modules are the same as RegT-GCN, but use random connected graphs as input subgraphs. RanT-GCN separates an original graph into a subgraph, which possesses the same number of nodes in each region with the graph decomposition, and each node is randomly selected and connected together. We also implemented the variants of RanT-GCN.
  • P: The total number of nodes in each region was randomly selected.
  • R: Each node was randomly selected to connect to each other.
  • S: Set the maximum degree of selected nodes to four.
Table IV: Different Connected Graph-Based Model Comparison at Time Horizon 10
RMSE MSE MAPE
RanT-GCN 0.090 0.066 9.124
RanT-GCN (R) 0.090 0.066 9.125
RanT-GCN (P) 0.089 0.065 9.080
RanT-GCN (S) 0.089 0.064 9.14
RegT-GCN 0.086 0.062 8.75
In the study of different connectivity approaches, RanT-GCN and RegT-GCN outperform the baseline models in most of the prediction time horizons. Within these two, regionally decomposing a graph achieved the best accuracy. Table IV explains that node selection contributes more than changing the degree of nodes and regional selection. However, the experiments on different node selections (RanT-GCN vs. RanT-GCN (R)) came to the same results; thus, the node selection should be taken carefully, as we will mention in our discussion.
IV-G Discussion